A botanist wishes to estimate the typical number of seeds for a certain fruit. She samples 42 specimens and counts the number of seeds in each. Use her sample results (mean = 79.1, standard deviation = 10.3) to find the 95% confidence interval for the number of seeds for the species. Enter your answer as an open-interval (i.e., parentheses) accurate to 3 decimal places.  

95% C.I. =
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 79.1

sample standard deviation = s = 10.3

sample size = n = 42

Degrees of freedom = df = n - 1 = 41

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,41 = 2.020

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.020* (10.3 / \sqrt 42)

= 3.210

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

79.1 - 3.210 < \mu < 79.1 + 3.210

75.890 < \mu < 82.310

95% C.I. = ( 75.890 , 82.310)
Answer #2
A botanist wishes to estimate the typical number of seeds for a certain fruit. She samples 67 specimens and counts the number of seeds in each. Use her sample results (mean =49.8, standard. deviation = 9.7) to find the 95% confidence interval for the number of seeds for the species. Enter your answer as an open- interval.
source: Statistics
answered by: Shamain